OpenStudy (anonymous):

how do you find the interception point of two lines using the coordinates (3,4) ,(11,-2) for line 1 and (9,6),(1,-2) for line two. I did the slope intercept form and got the coordinate (13/7,-8/7) but this point is no where near the point of interception when I graph the two lines

OpenStudy (anonymous):

Finding the slope of line 1 I get -3/4, of 2, 1. The point-slope equations are (-3/4)x+(25/4), and x-3. Setting these equal to each other and solving for x, I get x = 37/7, or 5.29, so the point is (37/7, 16/7).

OpenStudy (anonymous):

can you show me the work for solving for x so I can see what I did wrong/

OpenStudy (anonymous):

x-3 = (-3/4)x+(25/4) x+(3/4)x = (25/4)+3 (7/4)x = (37/4) 7x = 37 x = 37/7 Makes sense?

OpenStudy (anonymous):

Okay, I see where I wrong. Forgot a negative. Thanks!